OpenStudy (sleepyjess):

Expand the following using either the Binomial Theorem or Pascal’s Triangle. (2x + 4)^3

OpenStudy (sleepyjess):

@ParthKohli

OpenStudy (kirbykirby):

\[(2x+4)^3=\sum_{r=0}^3{ 3\choose r}(2x)^{3-r}(4)^r \\ = { 3\choose 0}(2x)^{3-0}4^0+{ 3\choose 1}(2x)^{3-1}(4)^1+{ 3\choose 2}(2x)^{3-2}4^2+{ 3\choose 3}(2x)^{3-3}(4)^3\]

OpenStudy (sleepyjess):

Is that what I have to do?

OpenStudy (kirbykirby):

That's using the binomial theorem, but yes. You can simplify the expression more by evaluating each binomial coefficient, since \[ {n \choose r}=\frac{n!}{r!(n-r)!}\] or by using Pascal's triangle

OpenStudy (sleepyjess):

Thank you so much! The lesson didn't really explain either so I had no clue of what I was supposed to do.

OpenStudy (kirbykirby):

ah well yw :)